Ask
your own question, for FREE!
Mathematics
9 Online
We are drawing two cards without replacement from a standard 52 card deck. Find the probability that we draw at least one queen.
Still Need Help?
Join the QuestionCove community and study together with friends!
the probability of drawing 1 queen in 1 pick is 1/13. Since you have two chances, you get 1/13 + 1/13 = 2/13
That's not right.
The probability of getting at least one queen is the probability of getting a queen on the first draw plus the probability of getting something else on the first draw and a queen on the second. So the probability is: 1/13+48/51 * 4/51
@Thomas9 the question requires AT LEAST ONE QUEEN, not exactly one queen.
I'm correct, you can get two queens in my method, it's in the 1/13 term. To disprove your method: if you draw 14 cards you'd end up with a probability of 14/13. which is absurd.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
18 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal